OpenStudy (anonymous):

I need help finding the solution set for the following inequality:6-5/4*x is greater than or equal to 9

OpenStudy (yttrium):

Hey. If I got what you mean, maybe this is the answer. 6-(5/4x)>= 9, Compute 1st for the LCD of the left side: (24x-5)/4x >= 9 , Then cross multiply: 24x-5 >= 36x, Combine similar terms 12x>=-5, Therefore x>= -5/12

OpenStudy (anonymous):

Is that the solution set? That is the answer I got also but I thought a solution set was in the form of, for example: 5<=x<=1.

OpenStudy (anonymous):

I actually got x<=-12/5

OpenStudy (anonymous):

Shouldn't the sign be flipped?

OpenStudy (yttrium):

it's already flipped, yeah that's already a solution set. It does not necesarily mean that a solution set must be something like x<= x <= 1, it's 2 directional well the answer here is just one directional solution set.

OpenStudy (anonymous):

But you got x>= and I got x<=

OpenStudy (yttrium):

Can you show me your solution, I already flipped the inequality sign since I've got the negative x well if you want, can you show me you solution so as to attest it. :D

OpenStudy (yttrium):

Uhh. I'm sorry, I lost my mind during axioms. You're correct. :D

OpenStudy (yttrium):

You're welcome. :D